# Distributionally Robust Policy Learning under Concept Drifts

### v1

@@ 2.2 Strong duality
In order to estimate $\mathcal{V}_{\delta}(\pi)$, we first rewrite the inner optimization problem in Equation (1) in its dual form using standard results in convex optimization (see e.g., Luenberger, (1997)). The transformation is formalized in the following lemma, with its proof provided in Appendix B.1.
@@
$-\min_{\alpha\geq 0,\eta\in\mathbb{R}}\mathbb{E}_{P}\bigg{[}\alpha\exp\Big{(}-\frac{Y(\pi(X))+\eta}{\alpha}-1\Big{)}+\eta+\alpha\delta\,\Big{|}\,X=x\bigg{]}.$ (3)
@@
$\big{(}\bm{\alpha}^{*}_{\pi}(x),\bm{\eta}^{*}_{\pi}(x)\big{)}\in\underset{\alpha\geq 0,\eta\in\mathbb{R}}{\arg\!\min}~{}\mathbb{E}_{P}\bigg{[}\alpha\exp\Big{(}{-\frac{Y(\pi(X))+\eta}{\alpha}-1}\Big{)}+\eta+\alpha\delta\,\Big{|}\,X=x\bigg{]}.$
With this notation and Lemma 2.3, the robust policy value becomes
$\mathcal{V}_{\delta}(\pi)=-\mathbb{E}_{P}\bigg{[}\bm{\alpha}^{*}_{\pi}(X)\exp\Big{(}-\frac{Y(\pi(X))+\bm{\eta}^{*}_{\pi}(X)}{\bm{\alpha}^{*}_{\pi}(X)}-1\Big{)}+\bm{\eta}^{*}_{\pi}(X)+\bm{\alpha}^{*}_{\pi}(X)\delta\bigg{]}.$ (4)
@@
We end this section by providing a sufficient condition to ensure $\bm{\alpha}^{*}_{\pi}(x)>0$, which we shall make throughout the paper.
@@
The above assumption is mild and can be satisfied by many commonly used distributions, e.g., all the continuous distributions; it requires that $P_{Y\,|\,X,A}$ does not posit a large point mass at its essential infimum. Next, the following result from Jin et al., 2022b (, Proposition 4), shows that $\alpha^{*}>0$ when Assumption 2.4 holds, which ensures that the gradient of the risk function in ERM has a zero mean.
Proposition 2.5 (Jin et al., 2022b ).
@@ 3.1 The estimation procedure
For each $k\in[K]$, we first use data points in $\mathcal{D}^{(k+1)}$ to obtain the propensity score estimator $\widehat{\pi}_{0}^{(k)}$ and the optimizers $(\widehat{\bm{\alpha}}_{\pi}^{(k)},\widehat{\bm{\eta}}_{\pi}^{(k)})$. [^4] We use the convention that $\mathcal{D}^{(k+j)}=\mathcal{D}^{(k+j\textnormal{ mod }K)}$ for any $j,k$. We then define
$\widehat{G}_{\pi}^{(k)}(x,y)$ $:=\widehat{\bm{\alpha}}_{\pi}^{(k)}(x)\cdot\exp\Big{(}-\frac{y+\widehat{\bm{\eta}}_{\pi}^{(k)}(x)}{\widehat{\bm{\alpha}}_{\pi}^{(k)}(x)}-1\Big{)}+\widehat{\bm{\eta}}_{\pi}^{(k)}(x)+\widehat{\bm{\alpha}}_{\pi}^{(k)}(x)\cdot\delta,$
and its conditional expectation $\bar{g}^{(k)}_{\pi}(x):=\mathbb{E}_{P}\Big{[}\widehat{G}_{\pi}^{(k)}\big{(}X,Y(\pi(X))\big{)}\,\big{|}\,X=x\Big{]}$. We next use $\mathcal{D}^{(k+2)}$ to obtain $\widehat{g}_{\pi}^{(k)}$ as an estimator of $g_{\pi}$. The policy value estimator $\widehat{\mathcal{V}}_{\delta}^{(k)}(\pi)$ for the $k$-th fold is constructed as
$\widehat{\mathcal{V}}_{\delta}^{(k)}(\pi)=\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i\in\mathcal{D}^{(k)}}\frac{\mathbbm{1}\{\pi(X_{i})=A_{i}\}}{\widehat{\pi}_{0}^{(k)}(A_{i}\,|\,X_{i})}\cdot\big{(}\widehat{G}_{\pi}^{(k)}(X_{i},Y_{i})-\widehat{g}_{\pi}^{(k)}(X_{i})\big{)}+\widehat{g}_{\pi}^{(k)}(X_{i}).$ (5)
The final policy value estimator is given by $\widehat{\mathcal{V}}_{\delta}(\pi):=-\frac{1}{K}\sum^{K}_{k=1}\widehat{\mathcal{V}}_{\delta}^{(k)}(\pi)$. The complete procedure is summarized in Algorithm 1.
@@
     On $\mathcal{D}^{(k+1)}$: $\widehat{\pi}_{0}^{(k)}\leftarrow\mathcal{C}(\mathcal{D}^{(k+1)})$, $(\widehat{\bm{\alpha}}^{(k)}_{\pi},\widehat{\bm{\eta}}_{\pi}^{(k)})\leftarrow\mathcal{E}(\mathcal{D}^{(k+1)})$;
     On $\mathcal{D}^{(k+2)}$: $\widehat{g}_{\pi}^{(k)}\leftarrow\mathcal{R}\big{(}\{X_{i},A_{i},\widehat{G}_{\pi}^{(k)}(X_{i},Y_{i});i\in\mathcal{D}^{(k+2)}\}\big{)}$;
     On $\mathcal{D}^{(k)}$: compute $\widehat{\mathcal{V}}_{\delta}^{(k)}(\pi)$ according to Equation (5);
@@
The estimation procedure involves three model-fitting steps corresponding to $\pi_{0}$, $(\bm{\alpha}^{*}_{\pi},\bm{\eta}^{*}_{\pi})$, and $\bar{g}_{\pi}$, respectively. The propensity score function $\pi_{0}$ can be estimated with off-the-shelf algorithms (e.g., logistic regression, random forest); the conditional mean $\bar{g}_{\pi}^{(k)}$ can be obtained by regressing $\widehat{G}_{\pi}^{(k)}(X_{i},Y_{i})$ onto $X_{i}$ for the points such that $A_{i}=\pi(X_{i})$ with standard regression algorithms, e.g., kernel regression (Nadaraya,, 1964; Watson,, 1964), local polynomial regression (Cleveland,, 1979; Cleveland and Devlin,, 1988), smoothing spline (Green and Silverman,, 1993), regression trees (Loh,, 2011) and random forests (Ho et al.,, 1995). The ERM step is more complex, and will be discussed in detail shortly.
@@
The construction of the estimator $\widehat{\mathcal{V}}_{\delta}(\pi)$ employs two major techniques: cross-fitting and de-biasing. The cross-fitting technique crucially provides the convenient property of independence and the de-biasing technique allows for the slow rate of estimating the nuisance parameter $\alpha_{\pi},\eta_{\pi}$, leading to the doubly-robust property of the proposed estimator.
The ERM step.
For notational simplicity, we denote $\theta=(\alpha,\eta)$ and write the loss function as
@@
For any policy $\pi$, there exist constants $\underline{\alpha},\bar{\alpha},\bar{\eta}$ such that
$0<\underline{\alpha}\leq\bm{\alpha}_{\pi}^{*}(x)\leq\bar{\alpha},\quad\big{|}\bm{\eta}_{\pi}^{*}(x)\big{|}\leq\bar{\eta},\quad\text{for all }x\in\mathcal{X}.$
The above assumption is mild and can be achieved, for example, when $\bm{\theta}^{*}_{\pi}(x)$ is continuous in $x$ and when $\mathcal{X}$ is compact. We refer the readers to Jin et al., 2022b for a more detailed discussion.
Under the unconfoundedness assumption, it can be seen that $\bm{\theta}^{*}_{\pi}$ is also a minimizer of $\mathbb{E}_{P}\big{[}\ell(X,Y;\bm{\theta}(X))\mathbbm{1}\{A=\pi(X)\}\big{]}$. We can thus estimate $\bm{\theta}^{*}_{\pi}$ by minimizing the empirical risk:
$\widehat{\bm{\theta}}_{\pi}^{(k)}\in\underset{\bm{\theta}\in\Theta}{\arg\!\min}~{}\Bigg{\{}\frac{1}{|\mathcal{D}^{(k+1)}|}\sum_{i\in D^{(k+1)}}\mathbbm{1}\{A_{i}=\pi(X_{i})\}\cdot\ell\big{(}X_{i},Y_{i};\bm{\theta}(X_{i})\big{)}\Bigg{\}},$ (7)
where $\Theta\subseteq\{(\bm{\alpha},\bm{\eta})\mid\bm{\alpha}:\mathcal{X}\mapsto\mathbb{R}_{\geq 0},~{}\bm{\eta}:\mathcal{X}\mapsto\mathbb{R}\}$ is to be determined. In our implementation, we follow Yadlowsky et al., (2022); Jin et al., 2022b ; Sahoo et al., (2022), and adopt the method of sieves (Geman and Hwang,, 1982) to solve (7). Specifically, we consider an increasing sequence $\Theta_{1}\subset\Theta_{2}\subset\cdots$ of spaces of smooth functions, and let $\Theta=\Theta_{n}$ in Equation (7). For example, $\Theta_{n}$ can be a class of polynomials, splines, or wavelets. It has been shown in Jin et al., 2022b (, Section 3.4) that under mild regularity conditions, $\widehat{\bm{\theta}}_{\pi}^{(k)}$ converges to $\bm{\theta}^{*}_{\pi}$ at a reasonably fast rate. For example, if $\mathcal{X}=\prod^{d}_{j=1}\mathcal{X}_{j}\subseteq\mathbb{R}^{d}$ for some compact intervals $\mathcal{X}_{j}$ and that $\bm{\theta}^{*}_{\pi}$ belongs to the Hölder class of $p$-smooth functions — with some other mild regularity conditions — then $\|\widehat{\bm{\theta}}_{\pi}^{(k)}-\bm{\theta}_{\pi}^{*}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}=O_{P}((\frac{\log n}{n})^{-p/(2p+d)})$ and $\|\widehat{\bm{\theta}}_{\pi}^{(k)}-\bm{\theta}_{\pi}^{*}\|_{L_{\infty}}=O_{P}((\frac{\log n}{n})^{-2p^{2}/(2p+d)^{2}})$. We refer the readers to Yadlowsky et al., (2018) and Jin et al., 2022b for more details.
@@ 3.2 Theoretical guarantees
For any policy $\pi$, assume that for each $k\in[K]$,
  1. (1)
    The estimators $\widehat{\pi}_{0}^{(k)}$ and $\widehat{g}_{\pi}^{(k)}$ satisfy
    $\|\widehat{\pi}_{0}^{(k)}-\pi_{0}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}=o_{P}(n^{-\gamma_{1}}),~{}\|\widehat{g}_{\pi}^{(k)}-\bar{g}_{\pi}^{(k)}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}=o_{P}(n^{-\gamma_{2}}),$
@@
  2. (2)
    The empirical risk optimizer $\widehat{\bm{\theta}}_{\pi}^{(k)}$ satisfies
    $\|\widehat{\bm{\theta}}_{\pi}^{(k)}-\bm{\theta}^{*}_{\pi}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}=o_{P}(n^{-\frac{1}{4}}),~{}\|\widehat{\bm{\theta}}_{\pi}^{(k)}-\bm{\theta}^{*}_{\pi}\|_{L_{\infty}}=o_{P}(1).$
Assumption 3.4 (1) requires either the propensity score $\pi_{0}$ or the conditional mean of $\widehat{G}_{\pi}^{(k)}(X,Y)$ is well estimated. This is a standard assumption in the double machine learning literature (Chernozhukov et al.,, 2018; Athey and Wager,, 2021; Zhou et al.,, 2023; Kallus et al.,, 2019, 2022; Jin et al., 2022b, ) and can be achieved by various commonly-used machine learning methods discussed in Section 3.1. Assumption 3.4 (2) requires the optimizer $\widehat{\bm{\theta}}_{\pi}^{(k)}$ to be estimated at a rate faster than $n^{-1/4}$, and can be achieved by, for example, the estimators discussed in Section 3.1 under mild conditions.
The following theorem states that our estimated policy value $\widehat{\mathcal{V}}_{\delta}(\pi)$ is consistent for estimating $\mathcal{V}_{\delta}$ and is asymptotically normal. Its proof is provided in Appendix B.2.
@@
Suppose Assumptions 2.12.43.3, and 3.4 hold. For any policy $\pi:\mathcal{X}\mapsto[M]$, we have
$\sqrt{n}\cdot\big{(}\widehat{\mathcal{V}}_{\delta}(\pi)-\mathcal{V}_{\delta}(\pi)\big{)}\stackrel{{\scriptstyle\textnormal{d}}}{{\rightarrow}}N(0,\sigma^{2}_{\pi}),$
where

### v2

@@ 2.2 Strong Duality
In order to estimate $\mathcal{V}_{\delta}(\pi)$, we first rewrite the inner optimization problem in Equation (1) in its dual form using standard results in convex optimization (see e.g., Luenberger (1997)). The transformation is formalized in the following lemma, with its proof provided in Appendix D.1.
@@
$-\min_{\alpha\geq 0,\eta\in\mathbb{R}}\mathbb{E}_{P}\bigg{[}V_{\delta}(\alpha,\eta;\pi)\,\Big{|}\,X=x\bigg{]}.$ (3)
where $V_{\delta}(\alpha,\eta;\pi)=\alpha\exp(-\frac{Y(\pi(X))+\eta}{\alpha}-1)+\eta+\alpha\delta$.
@@
$\big{(}\bm{\alpha}^{*}_{\pi}(x),\bm{\eta}^{*}_{\pi}(x)\big{)}\in\underset{\alpha\geq 0,\eta\in\mathbb{R}}{\arg\!\min}~{}\mathbb{E}_{P}\bigg{[}V_{\delta}(\alpha,\eta;\pi)\,\Big{|}\,X=x\bigg{]}.$
With this notation and Lemma 2.3, the robust policy value
$\mathcal{V}_{\delta}(\pi)=-\mathbb{E}_{P}\bigg{[}V_{\delta}(\bm{\alpha}^{*}_{\pi}(X),\bm{\eta}^{*}_{\pi}(X);\pi)\bigg{]}.$ (4)
@@
We end this section by providing a sufficient condition to ensure $\bm{\alpha}^{*}_{\pi}(x)>0$, which we make throughout the paper.
@@
The above assumption is mild and can be satisfied by many commonly used distributions, e.g., all the continuous distributions; it requires that $P_{Y\,|\,X,A}$ does not posit a large point mass at its essential infimum. The following result from Jin et al. (2022b, Proposition 4), shows that $\alpha^{*}>0$ when Assumption 2.4 holds, ensuring that the gradient of the risk function in ERM has a zero mean.
Proposition 2.5 (Jin et al. (2022b)).
@@ 3.1 The Estimation Procedure
For each $k\in[K]$, we first use data points in $\mathcal{D}^{(k+1)}$ to obtain the propensity score estimator $\widehat{\pi}_{0}^{(k)}$ and the optimizers $(\widehat{\bm{\alpha}}_{\pi}^{(k)},\widehat{\bm{\eta}}_{\pi}^{(k)})$. [^4] We use the convention that $\mathcal{D}^{(k)}=\mathcal{D}^{(k\textnormal{ mod }K)}$ for any $k$. We then define
$\widehat{G}_{\pi}^{(k)}(x,y):=$ $\widehat{\bm{\alpha}}_{\pi}^{(k)}(x)\cdot e^{-\frac{y+\widehat{\bm{\eta}}_{\pi}^{(k)}(x)}{\widehat{\bm{\alpha}}_{\pi}^{(k)}(x)}-1}+\widehat{\bm{\eta}}_{\pi}^{(k)}(x)+\widehat{\bm{\alpha}}_{\pi}^{(k)}(x)\cdot\delta,$
$\bar{g}^{(k)}_{\pi}(x):=$ $\mathbb{E}_{P}\Big{[}\widehat{G}_{\pi}^{(k)}\big{(}X,Y(\pi(X))\big{)}\,\big{|}\,X=x\Big{]}.$
We next use $\mathcal{D}^{(k+2)}$ to obtain $\widehat{g}_{\pi}^{(k)}$ as an estimator of $\bar{g}_{\pi}^{(k)}$. The policy value estimator $\widehat{\mathcal{V}}_{\delta}^{(k)}(\pi)$ for the $k$-th fold is
$\widehat{\mathcal{V}}_{\delta}^{(k)}(\pi)=$ $\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i\in\mathcal{D}^{(k)}}\frac{\mathbbm{1}\{\pi(X_{i})=A_{i}\}}{\widehat{\pi}_{0}^{(k)}(A_{i}\,|\,X_{i})}$
$\cdot\big{(}\widehat{G}_{\pi}^{(k)}(X_{i},Y_{i})-\widehat{g}_{\pi}^{(k)}(X_{i})\big{)}+\widehat{g}_{\pi}^{(k)}(X_{i}).$ (5)
The policy value estimator is $\widehat{\mathcal{V}}_{\delta}(\pi):=-\frac{1}{K}\sum^{K}_{k=1}\widehat{\mathcal{V}}_{\delta}^{(k)}(\pi)$. The complete procedure is summarized in Algorithm 1.
@@
     $\widehat{\pi}_{0}^{(k)}\leftarrow\mathcal{C}(\mathcal{D}^{(k+1)})$, $(\widehat{\bm{\alpha}}^{(k)}_{\pi},\widehat{\bm{\eta}}_{\pi}^{(k)})\leftarrow\mathcal{E}(\mathcal{D}^{(k+1)})$;
     $\widehat{g}_{\pi}^{(k)}\leftarrow\mathcal{R}\big{(}\{X_{i},A_{i},\widehat{G}_{\pi}^{(k)}(X_{i},Y_{i});i\in\mathcal{D}^{(k+2)}\}\big{)}$;
     Compute $\widehat{\mathcal{V}}_{\delta}^{(k)}(\pi)$ according to Equation (3.1);
@@
The estimation procedure involves three model-fitting steps corresponding to $\pi_{0}$, $(\bm{\alpha}^{*}_{\pi},\bm{\eta}^{*}_{\pi})$, and $\bar{g}_{\pi}$, respectively. The propensity score function $\pi_{0}$ can be estimated with off-the-shelf algorithms (e.g., logistic regression, random forest); the conditional mean $\bar{g}_{\pi}^{(k)}$ can be obtained by regressing $\widehat{G}_{\pi}^{(k)}(X_{i},Y_{i})$ onto $X_{i}$ for the points such that $A_{i}=\pi(X_{i})$ with standard regression algorithms, e.g., kernel regression (Nadaraya, 1964; Watson, 1964), local polynomial regression (Cleveland, 1979; Cleveland & Devlin, 1988), smoothing spline (Green & Silverman, 1993), regression trees (Loh, 2011) and random forests (Ho et al., 1995). The ERM step is more complex, and will be discussed in details shortly.
@@
The construction of the estimator $\widehat{\mathcal{V}}_{\delta}(\pi)$ employs two major techniques: cross-fitting and de-biasing. The cross-fitting technique splits the training dataset $\mathcal{D}$ into $K$ folds equally and fits the nuisance parameters on the off-fold data. This crucially provides the convenient property of independence between the fitted nuisance parameters and policy value estimators. In contrast to the naïve plug-in estimator whose convergence rate is largely affected by the slow estimation rate of the nuisance parameters, the de-biasing technique addresses this limitation, thereby leading to the doubly-robust property of the proposed estimator.
The ERM Step. For notational simplicity, we denote $\theta=(\alpha,\eta)$ and write the loss function from Lemma 2.3 as
@@
For any policy $\pi$, $\exists\underline{\alpha},\bar{\alpha},\bar{\eta}$ such that $0<\underline{\alpha}\leq\bm{\alpha}_{\pi}^{*}(x)\leq\bar{\alpha}$, $\big{|}\bm{\eta}_{\pi}^{*}(x)\big{|}\leq\bar{\eta}$, $\forall x\in\mathcal{X}$.
The above assumption is mild and can be achieved, for example, when $\bm{\theta}^{*}_{\pi}(x)$ is continuous in $x$ and when $\mathcal{X}$ is compact. We refer the readers to (Jin et al., 2022b) for a more detailed discussion.
Under the unconfoundedness assumption, $\bm{\theta}^{*}_{\pi}$ is also a minimizer of $\mathbb{E}_{P}\big{[}\ell(X,Y;\bm{\theta}(X))\mathbbm{1}\{A=\pi(X)\}\big{]}$. We can thus estimate $\bm{\theta}^{*}_{\pi}$ by minimizing the empirical risk:
$\widehat{\bm{\theta}}_{\pi}^{(k)}\in\underset{\bm{\theta}\in\Theta}{\arg\!\min}~{}\Big{\{}\widehat{\mathbb{E}}_{\mathcal{D}^{(k+1)}}[\ell(X,Y;\bm{\theta}(X))\mathbbm{1}\{A=\pi(X)\}]\Big{\}},$ (7)
where $\Theta\subseteq\{(\bm{\alpha},\bm{\eta})\mid\bm{\alpha}:\mathcal{X}\mapsto\mathbb{R}_{\geq 0},~{}\bm{\eta}:\mathcal{X}\mapsto\mathbb{R}\}$ is to be determined. In our implementation, we follow Yadlowsky et al. (2022); Jin et al. (2022b); Sahoo et al. (2022), and adopt the method of sieves (Geman & Hwang, 1982) to solve (7). Specifically, we consider an increasing sequence $\Theta_{1}\subset\Theta_{2}\subset\cdots$ of spaces of smooth functions, and let $\Theta=\Theta_{n}$ in Equation (7). For example, $\Theta_{n}$ can be a class of polynomials, splines, or wavelets. It has been shown in Jin et al. (2022b, Section 3.4) that under mild regularity conditions, $\widehat{\bm{\theta}}_{\pi}^{(k)}$ converges to $\bm{\theta}^{*}_{\pi}$ at a reasonably fast rate. For example, if $\mathcal{X}=\prod^{d}_{j=1}\mathcal{X}_{j}\subseteq\mathbb{R}^{d}$ for some compact intervals $\mathcal{X}_{j}$ and that $\bm{\theta}^{*}_{\pi}$ belongs to the Hölder class of $p$-smooth functions — with some other mild regularity conditions — then $\|\widehat{\bm{\theta}}_{\pi}^{(k)}-\bm{\theta}_{\pi}^{*}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}=O_{P}((\frac{\log n}{n})^{-p/(2p+d)})$ and $\|\widehat{\bm{\theta}}_{\pi}^{(k)}-\bm{\theta}_{\pi}^{*}\|_{L_{\infty}}=O_{P}((\frac{\log n}{n})^{-2p^{2}/(2p+d)^{2}})$. The solution details are given in Appendix B, and we also refer the readers to Yadlowsky et al. (2018) and Jin et al. (2022b).
@@ 3.2 Theoretical Guarantees
For any policy $\pi$, assume that for each $k\in[K]$, the estimators $\widehat{\pi}_{0}^{(k)}$, $\widehat{g}_{\pi}^{(k)}$ and the empirical risk optimizer $\widehat{\bm{\theta}}_{\pi}^{(k)}$ satisfy
$\|\widehat{\pi}_{0}^{(k)}-\pi_{0}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}$ $=o_{P}(n^{-\gamma_{1}}),$
$\|\widehat{g}_{\pi}^{(k)}-\bar{g}_{\pi}^{(k)}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}$ $=o_{P}(n^{-\gamma_{2}}),$
$\|\widehat{\bm{\theta}}_{\pi}^{(k)}-\bm{\theta}^{*}_{\pi}\|_{L_{2}(P_{X\,|\,A=\pi(X)})}$ $=o_{P}(n^{-\frac{1}{4}}),$
$\|\widehat{\bm{\theta}}_{\pi}^{(k)}-\bm{\theta}^{*}_{\pi}\|_{L_{\infty}}$ $=o_{P}(1),$
@@
Assumption 3.4 (1) requires either the propensity score $\pi_{0}$ or the conditional mean of $\widehat{G}_{\pi}^{(k)}(X,Y)$ is well estimated. This is a standard assumption in the double machine learning literature (Chernozhukov et al., 2018; Athey & Wager, 2021; Zhou et al., 2023; Kallus et al., 2019, 2022; Jin et al., 2022b) and can be achieved by various commonly-used machine learning methods discussed in Section 3.1. Assumption 3.4 (2) requires the optimizer $\widehat{\bm{\theta}}_{\pi}^{(k)}$ to be estimated at a rate faster than $n^{-1/4}$, and can be achieved by, for example, the estimators discussed in Section 3.1 under mild conditions. The empirical sensitivity analysis in Jin et al. (2022b) also provides some justification for Assumption 3.4.
The following theorem states that our estimated policy value $\widehat{\mathcal{V}}_{\delta}(\pi)$ is consistent for estimating $\mathcal{V}_{\delta}$ and is asymptotically normal. Its proof is provided in Appendix D.2.
@@
Suppose Assumptions 2.12.43.3, and 3.4 hold. For any policy $\pi:\mathcal{X}\mapsto[M]$, we have $\sqrt{n}\cdot\big{(}\widehat{\mathcal{V}}_{\delta}(\pi)-\mathcal{V}_{\delta}(\pi)\big{)}\stackrel{{\scriptstyle\textnormal{d}}}{{\rightarrow}}N(0,\sigma^{2}_{\pi})$, where